The sum of the ages of Sergio and Adriana is 114 years. 7 years ago, Sergio's age was 4 times Adriana's age. How old is Sergio now?
Found 2 solutions by josgarithmetic, ikleyn:
Answer by josgarithmetic(39618)

Similar to the Jake and Maria problem.

Answer by ikleyn(52797)

.


7 years ago, the sum of their ages was  114 - 2*7 = 100 years, and Sergio was 4 times Adriana's age.


Hence, 7 years ago Adriana was 100:5 = 20 years and Sergio was 4*20 = 80 years old.


Hence, now Adriana is 20+7 = 27 years old;  Sergio is 80+7 = 87 years old.    ANSWER

Solved.